A stained glass window has 25 blue pieces and 20 red pieces. If 2 pieces are selected at random, what is P(2 red or 2 blue)?
Answer by stanbon(57290)

A stained glass window has 25 blue pieces and 20 red pieces. If 2 pieces are selected at random, what is P(2 red or 2 blue)?
---
red and blue are mutually exclusive events.
----
P(2red OR 2blue) = P(2red) + P(2blue) = 20C2/45C2 + 25C2/45C2
------
= [20C2+25C2]/45C2 = 490/990 = 0.4949
==========================================
Cheers,
Stan H.
=========